# ResGS: Residual Densification of 3D Gaussian for Efficient Detail Recovery

### v2

@@ Abstract
Recently, 3D Gaussian Splatting (3D-GS) has prevailed in novel view synthesis, achieving high fidelity and efficiency. However, it often struggles to capture rich details and complete geometry.
Our analysis reveals that the 3D-GS densification operation lacks adaptiveness and faces a dilemma between geometry coverage and detail recovery. To address this, we introduce a novel densification operation, residual split, which adds a downscaled Gaussian as a residual. Our approach is capable of adaptively retrieving details and complementing missing geometry. To further support this method, we propose a pipeline named ResGS. Specifically, we integrate a Gaussian image pyramid for progressive supervision and implement a selection scheme that prioritizes the densification of coarse Gaussians over time. Extensive experiments demonstrate that our method achieves SOTA rendering quality. Consistent performance improvements can be achieved by applying our residual split on various 3D-GS variants, underscoring its versatility and potential for broader application in 3D-GS-based applications.
@@ 1 Introduction
Novel View Synthesis (NVS) has continued to be a hot topic throughout the years. Traditional methods like point clouds and meshes [4, 14, 9] can achieve high rendering speed but often lack fidelity. On the other hand, neural volumetric approaches [31, 1, 32] can achieve high fidelity for novel views but significantly sacrifice rendering efficiency. Recently, 3D-GS [24] has shown both high fidelity and fast rendering speed for novel view synthesis. However, 3D-GS still suffers from setbacks, often failing to capture fine details and recover complete geometry, as shown by red and orange rectangles in Fig. 1 and Fig. 2.
@@
While various approaches [48, 12, 50, 43, 13] have been proposed to tackle these challenges and have significantly advanced the rendering quality of 3D-GS-based methods to SOTA levels, they have not modified the densification operation of 3D-GS. In this paper, we take a novel approach by revisiting and refining the densification operation to improve reconstruction quality.
@@
The densification operation used in 3D-GS-based methods consists of two operations: split, designed to address over-reconstruction (where Gaussians cover too much area, resulting in a lack of detail), and clone, aimed at under-reconstruction (missing geometry). However, selecting between the two operations is challenging. 3D-GS-based methods rely on a threshold-based decision mechanism. As illustrated in Fig. 1 (a), Gaussians whose scale $s=max(s_{x},s_{y},s_{z})$ is larger than a predefined threshold $\tau_{s}$ undergo split, while those below it use clone. Due to its simplicity, the selected operation might be suboptimal and bring a dilemma in threshold selection. As highlighted by the red rectangles in Fig. 2 (a), a high threshold results in insufficient split, which prevents the model from capturing fine details, leading to blurry artifacts. On the other hand, a low threshold reduces clone operations, leading to missing geometry, as illustrated in the orange boxes in Fig. 2 (b). This dilemma limits the rendering quality. Furthermore, 3D-GS tends to first split and then clone, resulting in objects being fitted with Gaussians of similar scales. As highlighted by the blue rectangles in Fig. 1, 3D-GS fits textureless areas with an excessive number of small Gaussians, which is redundant.
@@
To tackle the aforementioned problems, we propose residual split, a novel densification operation. As illustrated in Fig. 1 (b), the core idea is that for any under-optimized Gaussian, we spawn a downscaled replica as a residual component while reducing the original Gaussian’s opacity. This simultaneously expands spatial coverage and injects finer details, eliminating the split/clone dilemma.
@@
However, directly applying residual split does not produce significant improvements. This is due to the inefficiency of initialization in 3D-GS. At early iterations, the scene is represented by a few coarse Gaussians that face both over-reconstruction and under-reconstruction. Gaussians added through densification are challenging to optimize, as they must address both issues simultaneously. It’s desired to develop an approach that decouples coverage and detail optimization.
@@
  • We develop a coarse-to-fine training pipeline, ResGS, equipped with residual split, which incorporates an image-pyramid-based supervision schedule and a Gaussian selection scheme to enhance rendering quality;
  • Our method achieves SOTA rendering quality in NVS on three datasets, including Mip-NeRF360 [1], Tanks&Temples [25], and DeepBlending [18]. Furthermore, applying residual split to multiple 3D-GS variants consistently brings performance improvements, demonstrating that our new densification operation is more effective than densification with a fixed scale threshold.
@@ 2.2 Progressive Training in NVS
Several progressive training schedules have been explored upon NeRF-based methods. BungeeNeRF [40] captured different levels of details in large city-scale datasets by gradually adding closer views to the training views and appending residual blocks to the MLP throughout the training process. Pyramid NeRF [51] proposed a progressive subdivision of the scene while utilizing an image pyramid to gradually add high-frequency information to the scene. BARF [28] proposed a scheduling approach to progressively incorporate higher frequency components in positional encoding, enabling the training of NeRFs without requiring camera poses.
@@ 3 Our Method
Figure 3: Overview of our ResGS pipeline. (a) The core of our pipeline, residual split, involves adding a downscaled replicate and then reducing the opacity of the original Gaussian. (b) We assign initial Gaussians a temporary attribute $l_{i}=0$ for densification selection, which is discarded after training. Next, the pipeline is split into $L$ ($L=3$) stages, with each stage trained on images downscaled using an image pyramid. Furthermore, we divide a single stage evenly into $K$ ($K=3$) substages for use in selecting Gaussians to densify. (c) The points selected for densification are determined by the substage $k$, $l_{i}$ and viewspace gradients of Gaussians. residual split is performed upon selected Gaussians, and $l_{j}$ would be assigned to new Gaussians.
@@ 3.1 Preliminaries
where a 3D vector $\mathbf{s}$ is used for the scaling matrix $\mathbf{S}$.
@@
$\alpha_{i}(\mathbf{x})=c_{i}G_{i}^{\prime}(\mathbf{x})$
@@
During the training process, 3D Gaussians are periodically densified. 3D-GS calculates the average viewspace gradients $\Delta L_{i}$ over a certain number (e.g. 100) of iterations to determine whether to perform densification. If $\Delta L_{i}$ exceeds threshold $\tau$, a split or clone action is triggered. The split operation addresses over-reconstruction, where excessive space is covered, while clone addresses under-reconstruction, where additional spatial coverage is needed. To determine whether to split or clone, 3DGS[24] sets a threshold $\tau_{s}=kR$, where $k$ is a predefined factor and $R$ is the scene extent. Split occurs when the scale $s=max(\mathbf{s}_{x},\mathbf{s}_{y},\mathbf{s}_{y})$ of a Gaussian surpasses a threshold $\tau_{s}$. Clone is applied when $s<\tau_{s}$.
@@ 3.2.1 Residual Split
The core idea of our densification operation is to generate downscaled Gaussians as residuals to supplement those requiring densification. The newly added Gaussian adaptively enhances the reconstructed model, allowing the scene to be refined as needed. The method is named as residual split. As illustrated in Fig. 3 (a), for a to-be-densified Gaussian $G_{i}$, our densification operation consists of two steps. Firstly, we add a down-scaled replicate $G_{j}$. Specifically, its opacity $o_{j}$, rotation $\mathbf{R}_{j}$, spherical harmonics coefficients $SH_{j}$ would be same as $G_{i}$, $\boldsymbol{\mu}_{j}$ is sampled using the original 3D Gaussian as a PDF for sampling, and $\mathbf{S}_{j}$ is defined as $\mathbf{S}_{i}$ divided by a set factor $\lambda_{s}$:
$\mathbf{S}_{j}=\frac{1}{\lambda_{s}}\mathbf{S}_{i}$ (4)
$\mathbf{R}_{j}=\mathbf{R}_{i},\ SH_{j}=SH_{i},\ o_{j}=o_{i}$
$\boldsymbol{\mu}_{j}\sim\mathcal{N}(\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})$
where $\boldsymbol{\Sigma}_{i}$ is the covariance matrix of $G_{i}$. Then, to tackle the density increase in the overlap of the two Gaussians, we reduce the opacity of the original Gaussian. This is accomplished by multiplying its opacity by a specified factor $\beta$:
$o^{\prime}_{i}=\beta o_{i}$ (5)
Figure 4: Intermediate results after applying residual split. We saved the Gaussians at 3,000 and 6,000 iterations after applying residual split. As observed, residual split can adaptively handle over-reconstruction and under-reconstruction by refining blurry regions with finer Gaussians while also effectively reconstructing missing structures. Zoom in for better view.
Unlike clone and split which lack adaptiveness, residual split can fill in missing geometry and recover details adaptively without requiring a trade-off, as shown in Fig. 2 (c). We further illustrate this using intermediate results, as shown in Fig. 4. In areas where details are challenging to recover, down-scaled Gaussians of finer granularity are progressively introduced to refine and reveal intricate details. On the other hand, since densification does not change the scale of the original Gaussian, the total coverage is increased after densification, helping to reconstruct missing structures. Moreover, our method’s nonidentical approach with hierarchical Gaussians of varying scales enables efficient shape fitting. As illustrated by the blue rectangles in Fig. 1, this hierarchy allows us to fill textureless regions with a few large Gaussians, significantly reducing redundancy. Consequently, our approach delivers higher rendering quality while using fewer Gaussians.
@@ 3.2.2 Image Pyramid as Supervision
Building on our motivation in Sec. 1, we designed a coarse-to-fine supervision schedule that gradually integrates information into the scene. Specifically, based on the original multi-view training pictures, we build an image pyramid $\{IP_{i}\}_{i=1}^{L}$ consisting of $L$ layers. $IP_{i}$ corresponds to the set of images at different viewpoints in layer $i$. $IP_{L}$ is the set of original pictures:
$IP_{i}=\{P^{v}_{i}\}_{v\in V}$ (6)
where $V$ is the set of viewpoints and $P^{v}_{i}$ refers to the picture at viewpoint $v$ in the $i^{\text{th}}$ layer. Let $(H_{L}^{v},W_{L}^{v})$ be the resolution of $P_{L}^{v}$, and $(H_{i}^{v},W_{i}^{v})$ represent the resolution of $P_{i}^{v}$. Then there is:
$(H_{i}^{v},W_{i}^{v})=(H_{L}^{v}/2^{L-i},W_{L}^{v}/2^{L-i})$ (7)
Correspondingly, our training process is divided into $L$ stages. At stage $i$, we train our model under the supervision of $IP_{i}$, increasing the resolution of the training pictures throughout the stages. Due to the variation of frequency bandwidth in different levels in the image pyramid, high-frequency information is gradually added to the scene. Thus, we can force our model to learn coarse and low-frequency features at early stages and construct an overall structure of the whole scene. At later stages, the focus shifts to capturing high-frequency components, carving details above the coarse expression.
@@ 3.2.3 Varying Gradient Threshold
To further emphasize progressive training, an incremental approach for selecting Gaussians to densify is designed. Specifically, as iterations progress, the model’s focus gradually shifts toward finer details. To support this, we aim to encourage the densification of coarse Gaussians to capture more detail over time. Due to the nonidentical of our residual split, we can expect the generated Gaussians to be finer than the original ones. This allows us to track the relative fineness level of Gaussians, denoted as $l_{i}$. We assign $l_{i}$ of initial Gaussians to 0. If the level of the Gaussian being densified is $l_{i}$, we assign the newly added Gaussian a level of $l_{i}+1$. Thus, as $l_{i}$ gets higher, Gaussians tend to be smaller and finer. Note that $l_{i}$ is only utilized during training and discarded afterward.
A single stage is then divided evenly into $K$ substages, creating $L\times K$ substages in total. The gradually increasing substage can be seen as an indicator of the scene details level. At substage $k$, we promote additional densification for Gaussians with $l_{i}<k$ and prioritize those at lower levels, achieved by applying varying gradient thresholds. Specifically, the threshold $\tau_{k,i}$ for a Gaussian $G_{i}$ is set as:
$\tau_{k,i}=\left\{\begin{aligned} &\tau,l_{i}\geq k\\ &\frac{\tau}{\alpha^{k-l_{i}}},l_{i}<k\end{aligned}\right.$ (8)
where $\tau$ is a set threshold and $\alpha>1$ is a chosen value.
4 Experiments
@@
4.1 Experiment Settings
Following previous works [24, 43, 48, 29, 35, 13], evaluation of our model was done on three datasets, including all nine scenes from Mip-NeRF360 [1], two scenes from Tanks&Temples [25], and two scenes from DeepBlending [18]. The test and train splits follow those of the original 3D-GS [24], and the image resolution is also maintained as in 3D-GS.
Three commonly used metrics were used to quantify the fidelity of rendered images: PSNR, SSIM [38], and LPIPS [49]. We also report the memory consumption as an indicator of the number of Gaussians.
We trained three versions of our model: two with AbsGS [43], leveraging its enhanced indication of Gaussians for densification—one standard version and one small version with higher gradient thresholds—and a third version using 3D-GS as the baseline. We applied the periodic opacity reduction from AbsGS across all three versions, as it proved to be a more effective alternative to the opacity reset used in 3D-GS. Additionally, we extended this reduction beyond the densification phase, finding it further minimized redundancy without compromising fidelity. Following all 3D-GS-like methods, we trained our model for 30k iterations. The densification stops at 12000 iterations. For the version with AbsGS, $\tau$ was set to $0.00067$, the small version to $0.0016$; for the version with 3D-GS, $\tau$ was set to $0.00028$. $L$ was set to $3$ and $K$ was set to $3$. Our first stage lasts for 2500 iterations, the second stage for 3500 iterations, and the rest iterations are in the last stage. A 500-iteration warm-up phase was introduced between stages with densification disabled. $\alpha$ was set to $2^{\frac{1}{3}}$, $\lambda_{s}$ was set to $1.6$ and $\beta$ was set to $0.3$. The loss function remains the same as in 3D-GS [24], consisting of a combination of $\mathcal{L}_{1}$ and D-SSIM losses. All experiments were done on an NVIDIA RTX3090 GPU.

### v3

@@ Abstract
Recently, 3D Gaussian Splatting (3D-GS) has prevailed in novel view synthesis, achieving high fidelity and efficiency. However, it often struggles to capture rich details and complete geometry. Our analysis reveals that the 3D-GS densification operation lacks adaptiveness and faces a dilemma between geometry coverage and detail recovery. To address this, we introduce a novel densification operation, residual split, which adds a downscaled Gaussian as a residual. Our approach is capable of adaptively retrieving details and complementing missing geometry. To further support this method, we propose a pipeline named ResGS. Specifically, we integrate a Gaussian image pyramid for progressive supervision and implement a selection scheme that prioritizes the densification of coarse Gaussians over time. Extensive experiments demonstrate that our method achieves SOTA rendering quality. Consistent performance improvements can be achieved by applying our residual split on various 3D-GS variants, underscoring its versatility and potential for broader application in 3D-GS-based applications. Project page: https://yanzhelyu.github.io/resgs.github.io/.
@@ 1 Introduction
Novel View Synthesis (NVS) has continued to be a hot topic throughout the years. Traditional methods like point clouds and meshes [4, 13, 9] can achieve high rendering speed but often lack fidelity. On the other hand, neural volumetric approaches [29, 1, 30] can achieve high fidelity for novel views but significantly sacrifice rendering efficiency. Recently, 3D-GS [23] has shown both high fidelity and fast rendering speed for novel view synthesis. However, 3D-GS still suffers from setbacks, often failing to capture fine details or recover complete geometry, as shown by the red and orange rectangles in Fig. 1 and Fig. 2.
@@
While various approaches [48, 50, 43, 12] have been proposed to tackle these challenges and have significantly advanced the rendering quality of 3D-GS-based methods to state-of-the-art (SOTA) levels, they seldom adapt the densification operation of 3D-GS. In this paper, we go through the densification procedure and propose a novel adaptive densification approach to improve reconstruction quality.
@@
The densification operation used in 3D-GS-based methods consists of two operations: split and clone. The split operation is designed to address over-reconstruction, where a single Gaussian covers too much area, and the clone operation aims to compensate for under-reconstruction. However, selecting between the two operations properly in different areas is challenging and existing methods rely on a threshold-based decision mechanism. As illustrated in Fig. 1 (a), Gaussians whose scale $s=max(s_{x},s_{y},s_{z})$ is larger than a predefined threshold $\tau_{s}$ undergo split, while those below it use clone. The simplicity of this threshold-based densification usually leads to suboptimal results and encounters a dilemma in threshold selection. As highlighted by the red rectangles in Fig. 2 (a), a high threshold results in insufficient split preventing the model from capturing fine details and leading to blurry rendering. On the other hand, a low threshold suppresses clone operations, leading to missing geometry, as illustrated in the orange boxes in Fig. 2 (b). Furthermore, 3D-GS tends to first split and then clone, resulting in objects being fitted with Gaussians of similar scales. As highlighted by the blue rectangles in Fig. 1, 3D-GS fits textureless areas with an excessive number of small Gaussians, which is redundant.
@@
To tackle the aforementioned problems, we propose residual split, a novel densification operation in 3D Gaussian splatting. As illustrated in Fig. 1 (b), the core idea is to spawn a downscaled replica as a residual component while reducing the original Gaussian’s opacity for any under-optimized Gaussian. This simultaneously expands spatial coverage and injects finer details, eliminating the dilemma of selecting split or clone.
@@
However, directly applying residual split does not produce significant improvements. This is due to the inefficiency the 3D-GS initialization. At early iterations, the scene is represented by a few coarse Gaussians that very likely face both over-reconstruction and under-reconstruction. Gaussians added through densification are challenging to optimize, as they must address both issues simultaneously. It’s desired to develop an approach that decouples coverage and detail optimization.
@@
  • We develop a coarse-to-fine training pipeline, ResGS. Equipped with the proposed residual split, it incorporates a multi-scale supervision with image pyramids and a Gaussian selection scheme to enhance rendering quality;
  • Our method achieves SOTA rendering quality in NVS on three datasets, including Mip-NeRF360 [1], Tanks&Temples [24], and DeepBlending [17]. Furthermore, applying our proposed residual split to multiple 3D-GS variants consistently brings performance improvements, demonstrating that our new densification operation is more effective than densification with a fixed scale threshold.
@@ 2.2 Progressive Training in NVS
Several progressive training schedules have been explored upon NeRF-based methods. BungeeNeRF [40] captured different levels of details in large city-scale datasets by gradually adding closer views to the training views and appending residual blocks to the MLP throughout the training process. Pyramid NeRF [51] employs a coarse-to-fine strategy, using an image pyramid to progressively introduce high-frequency details while simultaneously subdividing the scene representation.
@@ 3 Our Method
Figure 3: Overview of our ResGS pipeline. (a) The core of our pipeline, residual split, involves adding a downscaled replicate and then reducing the opacity of the original Gaussian. (b) We assign initial Gaussians a temporary attribute $l_{i}=0$ for densification selection, which is discarded after training. Next, the pipeline is split into $L$ stages, with each stage trained on images downscaled using an image pyramid. Each single stage is further divided evenly into $K$ substages, for selecting Gaussians to densify. (c) The points selected for densification are determined by the substage $k$, $l_{i}$, and viewspace gradients of Gaussians. Residual split is performed upon selected Gaussians, and $l_{j}$ would be assigned to new Gaussians.
@@ 3.1 Preliminaries
where a 3D vector $\mathbf{s}=({s}_{x},{s}_{y},{s}_{y})$ is used for the scaling matrix $\mathbf{S}$.
@@
$\alpha_{i}(\mathbf{x})=c_{i}G_{i}^{\prime}(\mathbf{x}),$
@@
During the training process, 3D Gaussians are periodically densified. 3D-GS calculates the average viewspace gradients $\Delta L_{i}$ over a certain number (e.g. 100) of iterations to determine whether to perform densification. If $\Delta L_{i}$ exceeds threshold $\tau$, a split or clone action is triggered. The split operation addresses over-reconstruction, where excessive space is covered, while clone addresses under-reconstruction, where additional spatial coverage is needed. To determine whether to split or clone, 3D-GS [23] sets a threshold $\tau_{s}=kR$, where $k$ is a predefined factor and $R$ is the scene extent. Split occurs when the scale $s=max({s}_{x},{s}_{y},{s}_{y})$ of a Gaussian surpasses the threshold $\tau_{s}$. Clone is applied when $s<\tau_{s}$.
@@ 3.2.1 Residual Split
The core idea of our densification operation is to generate downscaled Gaussians as residuals to supplement those requiring densification. The newly added Gaussian adaptively enhances the reconstructed model, allowing the scene to be refined as needed. The method is named as residual split. As illustrated in Fig. 3 (a), for a to-be-densified Gaussian $G_{i}$, our densification operation consists of two steps. Firstly, we add a down-scaled replicate $G_{j}$. Specifically, its opacity $o_{j}$, rotation $\mathbf{R}_{j}$, spherical harmonics coefficients $SH_{j}$ would be same as $G_{i}$, $\boldsymbol{\mu}_{j}$ is sampled using the original 3D Gaussian as a PDF for sampling, and $\mathbf{S}_{j}$ is defined as $\mathbf{S}_{i}$ divided by a predefined factor $\lambda_{s}$:
$\mathbf{S}_{j}=\frac{1}{\lambda_{s}}\mathbf{S}_{i},$ (4)
$\mathbf{R}_{j}=\mathbf{R}_{i},\ SH_{j}=SH_{i},\ o_{j}=o_{i},$
$\boldsymbol{\mu}_{j}\sim\mathcal{N}(\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i}),$
where $\boldsymbol{\Sigma}_{i}$ is the covariance matrix of $G_{i}$. Then, to tackle the density increase in the overlap of the two Gaussians, we reduce the opacity of the original Gaussian. This is accomplished by multiplying its opacity by a predefined factor $\beta$:
$o^{\prime}_{i}=\beta o_{i}.$ (5)
Figure 4: Intermediate results after applying residual split. We saved the Gaussians at 3,000 and 6,000 iterations after applying residual split. As observed, residual split can adaptively handle over-reconstruction and under-reconstruction by refining blurry regions with finer Gaussians while also effectively reconstructing missing structures.
While the clone and split in the original 3D-GS pipeline lack adaptiveness, our residual split can fill in missing geometry and recover details adaptively without requiring a trade-off, as shown in Fig. 2 (c). We further illustrate this using intermediate results, as shown in Fig. 4. In areas where details are challenging to recover, down-scaled Gaussians of finer granularity are progressively introduced to refine and reveal intricate details. On the other hand, since densification does not change the scale of the original Gaussian, the total coverage is increased after densification, helping to reconstruct missing structures. Moreover, our method’s approach to generate nonidentical Gaussians of varying scales enables efficient shape fitting. As illustrated by the blue rectangles in Fig. 1, this non-identicalness allows us to fill textureless regions with a few large Gaussians, significantly reducing redundancy. Consequently, our approach delivers higher rendering quality while using fewer Gaussians.
@@ 3.2.2 Image Pyramid as Supervision
To decouple coverage and detail optimization, we design a coarse-to-fine supervision schedule to gradually integrate fine high-frequency information into the scene. Specifically, based on the original multi-view training pictures, we build an image pyramid $\{\mathcal{I}_{i}\}_{i=1}^{L}$ consisting of $L$ layers. $\mathcal{I}_{i}$ corresponds to the set of images at different viewpoints in layer $i$. $\mathcal{I}_{L}$ is the set of original pictures:
$\mathcal{I}_{i}=\{P^{v}_{i}\}_{v\in V},$ (6)
where $V$ is the set of viewpoints and $P^{v}_{i}$ refers to the image at viewpoint $v$ in the $i^{\text{th}}$ layer. Let $(H_{L}^{v},W_{L}^{v})$ be the resolution of $P_{L}^{v}$, and $(H_{i}^{v},W_{i}^{v})$ represent the resolution of $P_{i}^{v}$. Then there is:
$(H_{i}^{v},W_{i}^{v})=(H_{L}^{v}/2^{L-i},W_{L}^{v}/2^{L-i}).$ (7)
Correspondingly, our training process is divided into $L$ stages. At the $i$-th stage, the Gaussians are trained under the supervision of $\mathcal{I}_{i}$. At the initial stage with lower image resolution, the 3D Gaussians are optimized to construct an overall structure of the whole scene from coarse and low-frequency features. As the image resolution increases, high-frequency details of the scene gradually emerge, shifting the focus toward adding finer, more intricate features.
@@ 3.2.3 Varying Gradient Threshold
To further emphasize the progressive training, we modify the strategy for selecting Gaussians to densify. As the training proceeds, the training focus is desired to shift toward finer details. Thus, we encourage the densification of coarse Gaussians to capture more detail over time. In our residual split, the generated Gaussians are finer than the original ones. Thus, we can track the relative fineness level of each Gaussian $G_{i}$, denoted as $l_{i}$. We assign $l_{i}$ of all initial Gaussians to 0. If the level of the Gaussian being densified is $l_{i}$, we assign the newly added Gaussian a level of $l_{i}+1$. Thus, as $l_{i}$ gets higher, Gaussians tend to be smaller and finer. Note that $l_{i}$ is only utilized during rendering and discarded after training.
A single stage is then divided evenly into $K$ substages, creating $L\times K$ substages in total. The gradually increasing substage can be seen as an indicator of the scene details level. At substage $k$, we promote additional densification for Gaussians with $l_{i}<k$ and prioritize those at lower levels, achieved by applying varying gradient thresholds. Specifically, the threshold $\tau_{k,i}$ for deciding whether densification is needed for a Gaussian $G_{i}$ is set as:
$\tau_{k,i}=\left\{\begin{aligned} &\tau,l_{i}\geq k,\\ &\frac{\tau}{\alpha^{k-l_{i}}},l_{i}<k,\end{aligned}\right$. (8)
where $\tau$ and $\alpha>1$ are predefined values.
4 Experiments and Results
@@
4.1 Experimental Settings
Following previous works [23, 43, 48, 27, 33, 12], evaluation of our model was done on three datasets, including all nine scenes from Mip-NeRF360 [1], two scenes from Tanks&Temples [24], and two scenes from Deep Blending [17]. The test and train splits follow those of the original 3D-GS [23], and the image resolution is also maintained as in 3D-GS. Three commonly used metrics were used to quantify the fidelity of rendered images: PSNR, SSIM [38], and LPIPS [49]. We also report the memory consumption as an indicator of the number of Gaussians.
We trained three versions of our model: two with AbsGS [43], leveraging its enhanced indication of Gaussians for densification—one standard version and one small version with higher gradient thresholds—and a third version using 3D-GS as the baseline. We applied the periodic opacity reduction from AbsGS across all three versions, as it proved to be a more effective alternative to the opacity reset used in 3D-GS. Additionally, we extended this reduction beyond the densification phase, finding it further minimized redundancy without compromising fidelity. Following previous methods, we trained our model for 30k iterations. The densification stops at 12000 iterations. For the version with AbsGS, $\tau=0.00067$, and $0.0016$ for the small version. For the version with 3D-GS, $\tau=0.00028$. We set $L=3$ and $K=3$. Our first stage lasts for 2500 iterations, the second stage for 3500 iterations, and the rest iterations are in the last stage. $\alpha=2^{\frac{1}{3}}$, $\lambda_{s}=1.6$ and $\beta=0.3$. The loss function remains the same as in 3D-GS [23], consisting of a combination of $\mathcal{L}_{1}$ and D-SSIM losses. All experiments were done on an NVIDIA RTX3090 GPU.